# On Predictability of Reinforcement Learning Dynamics for Large Language Models

### v2

@@ Appendix A Experimental Setup
We begin by outlining our experimental setup and relevant definitions. Let $\theta_{\text{init}}$ denote the parameters of a pretrained base LLM. By applying a training method $M$, we obtain the updated parameters $\theta_{\text{full}}$. In our experiments, we consider the following methods: Distillation (DIST) (Hinton et al., 2015), Supervised Fine-Tuning (SFT) (Ye et al., 2025; Wu et al., 2025), PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024; DeepSeek-AI et al., 2025), GRPO (DeepSeek-AI et al., 2025), Dr.GRPO (Liu et al., 2025) and DAPO (Yu et al., 2025).
@@
For DIST, we adopt the distilled model DeepSeek-R1-Distill-Qwen-7B and its base model Qwen2.5-Math-7B. For SFT, we adopt Qwen3-8B-Base as the base model, trained on the DeepMath-103K dataset with the LlamaFactory [^1] https://github.com/hiyouga/LLaMA-Factory framework. For PPO, we adopt the open-sourced Open-Reasoner-Zero-7B and its base model Qwen2.5-7B. For Dr.GRPO, we adopt the open-sourced Qwen2.5-Math-7B-Oat-Zero and its base model Qwen2.5-Math-7B. For RLOO and GRPO, we adopt Qwen3-8B-Base as the base model, trained on the DAPO-Math-17k dataset with the Verl [^2] https://github.com/volcengine/verl framework. For DAPO, we consider two variants: (i) ours (8B) — starting from Qwen3-8B-Base, trained on DAPO-Math-17k with Verl; (ii) open-source (32B)DAPO-Qwen-32B trained from the base Qwen2.5-32B with DAPO-Math-17k. All of our training runs are conducted on $8\times$ H800 80GB GPUs until the reward/loss converges.
@@
For Supervised Fine-Tuning (SFT), we adapt our training codebase with LlamaFactory (Sheng et al., 2025). We employ full-parameter training in Float16 precision, with the maximum sequence length set to 20,480 tokens. The training batch size is 1,024 and the mini-batch size is 8, corresponding to 256 gradient accumulation steps. The learning rate is set to $1\times 10^{-5}$ with warmup, and gradient clipping of 1.0 is applied. We monitor the training loss and terminate training once the loss decreases by less than $2\times 10^{-1}$ over five consecutive steps. We conduct the SFT training experiment on Qwen3-8B-Base models, using the DeepMath-107K (He et al., 2025) dataset. The chat template for SFT is specified as:
@@
For RLOO, GRPO, and DAPO, we adapt our training codebase with the Verl (Sheng et al., 2025) and follow the corresponding training setups. All methods share the same core configuration: the maximum prompt length is 2,048 tokens and the maximum response length is 20,480 tokens, yielding a total budget of 22,528 tokens. We adopt a training batch size of 512 and a mini-batch size of 32, corresponding to 16 gradient accumulation steps in Float16 precision. Each prompt generates n=8n{=}8 outputs during rollout. The learning rate is set to $1\times 10^{-6}$ with warmup, and gradient clipping of 1.0 is applied. We monitor the average reward per training batch and terminate training once the reward fails to improve for five consecutive steps.
@@
Appendix B In-depth Analysis of the Low-rank Phenomenon
@@
Second, common stabilization mechanisms—such as KL regularization, logits clipping, and gradient clipping—further constrain the magnitude and spread of parameter updates, thereby limiting, to some extent, the enrichment of update information. Importantly, our norm-based analysis (Figure 3(b)) demonstrates that even under such strong constraints, RL achieves substantial improvements in reasoning performance through limited updates. This suggests that performance gains do not rely on large-scale parameter drift but emerge from focused adjustments within a small set of critical subspaces.
@@
Appendix C External Manifestations of Rank-1 Dominance
@@
First, let’s make sure I understand all the components: Triangle ABCABC: A triangle with vertices AA, $B$, and $C$. Medians ADAD, BEBE, and CFCF: A median of a triangle is a line segment joining a vertex to the midpoint of the opposite side. So: $D$ is the midpoint of $BC$. $E$ is the midpoint of $AC$. $F$ is the midpoint of $AB$. Centroid $G$: The point where the three medians intersect. It’s also the “center of mass” of the triangle. Line through $G$ parallel to BCBC: This line is parallel to side BCBC and passes through the centroid $G$. Points $M$ and NN: The intersections of this parallel line with sides ABAB and $AC$, respectively. Area of triangle ABCABC: Given as 144. Goal: Find the area of triangle $ENG$. Note that point $E$ is already defined as the midpoint of $AC$, so triangle ENGENG has vertices $E$, NN, and $G$.
@@
Appendix D Linear Projection Methods

### v3

@@ Appendix B Experimental Setup
We begin by outlining our experimental setup and relevant definitions. Let $\theta_{\text{init}}$ denote the parameters of a pretrained base LLM. By applying a training method $M$, we obtain the updated parameters $\theta_{\text{full}}$.
In our experiments, we consider the following methods: Distillation (DIST) (Hinton et al., 2015), Supervised Fine-Tuning (SFT) (Ye et al., 2025; Wu et al., 2025), PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024; DeepSeek-AI et al., 2025), GRPO (DeepSeek-AI et al., 2025), Dr.GRPO (Liu et al., 2025b), DAPO (Yu et al., 2025), On-Policy Distillation (Agarwal et al., 2024), DPO (Rafailov et al., 2024), and Spiral (Liu et al., 2025a).
@@
For DIST, we adopt the distilled model DeepSeek-R1-Distill-Qwen-7B and its base model Qwen2.5-Math-7B. For SFT, we adopt Qwen3-8B-Base as the base model, trained on the DeepMath-103K dataset with the LlamaFactory [^1] https://github.com/hiyouga/LLaMA-Factory framework. For PPO, we adopt the open-sourced Open-Reasoner-Zero-7B and its base model Qwen2.5-7B. For Dr.GRPO, we adopt the open-sourced Qwen2.5-Math-7B-Oat-Zero and its base model Qwen2.5-Math-7B. For RLOO and GRPO, we adopt Qwen3-8B-Base as the base model, trained on the DAPO-Math-17k dataset with the Verl [^2] https://github.com/volcengine/verl framework. For DAPO, we evaluate the following models from small to large:
(1) 7BDeepSeek-R1-Distill-Qwen-7B, which is a distilled version of Qwen2.5-Math-7B , upon which we further trained on DAPO-Math-17k with Verl.
(2) 8B — starting from Qwen3-8B-Base, trained on DAPO-Math-17k with Verl.
(3) We perform cold-start training on Llama3.1-8B-Instruct with the DeepMath-103K dataset as the initialization, and subsequently perform DAPO with Verl.
(3) 9B — we additionally evaluate GLM-9B, trained on DAPO-Math-17k with Verl.
(4) 14B — we further evaluate Qwen3-14B-Base, trained on DAPO-Math-17k with Verl.
(5) 32BDAPO-Qwen-32B trained from the base Qwen2.5-32B using DAPO-Math-17k.
For On-Policy Distillation, We use Qwen3-8B-Base-Open-Thoughts-On-Policy-Distillation, whose base model is Qwen3-8B-Base.
For DPO, We use Qwen3-8B-Base, trained on Math-Step-DPO-10K with Verl.
For Spiral, we follow Liu et al. (2025a), and use Qwen3-4B-Base to perform adversarial training on Kuhn Poker and Simple Negotiation games.
All of our training runs are conducted on $8\times$ H800 80GB or $16\times$ H800 80GB GPUs until the reward/loss converges.
@@
For Supervised Fine-Tuning (SFT), we adapt our training codebase with LlamaFactory (Sheng et al., 2025). We employ full-parameter training in Float16 precision, with the maximum sequence length set to 20,480 tokens. The training batch size is 1,024 and the mini-batch size is 4, corresponding to 512 gradient accumulation steps. The learning rate is set to $1\times 10^{-5}$ with warmup, and gradient clipping of 1.0 is applied. We monitor the training loss and terminate training once the loss decreases by less than $2\times 10^{-1}$ over five consecutive steps. We conduct the SFT training experiment on Qwen3-8B-Base models, using the DeepMath-107K (He et al., 2025) dataset. The chat template for SFT is specified as:
@@
For RLOO, GRPO, and DAPO, we adapt our training codebase with the Verl (Sheng et al., 2025) and follow the corresponding training setups. All methods share the same core configuration: the maximum prompt length is 2,048 tokens and the maximum response length is 20,480 tokens, yielding a total budget of 22,528 tokens. During training, each backward pass uses a mini-batch of 32 samples, and the gradients are accumulated for 16 iterations before a single optimization step is performed, resulting in an effective batch size of 512 under Float16 precision. Each prompt generates n=8n{=}8 outputs during rollout. The learning rate is set to $1\times 10^{-6}$ with warmup, and gradient clipping of 1.0 is applied. We monitor the average reward per training batch and terminate training once the reward fails to improve for five consecutive steps.
@@
Appendix C In-depth Analysis of the Low-rank Phenomenon
@@
Second, common stabilization mechanisms—such as KL regularization, logits clipping, and gradient clipping—further constrain the magnitude and spread of parameter updates, thereby limiting, to some extent, the enrichment of update information. Importantly, our norm-based analysis (Figure 3(b)) demonstrates that even under such strong constraints, RL achieves substantial improvements in reasoning performance through limited updates. This suggests that performance gains do not rely on large-scale parameter drift but emerge from focused adjustments within a small set of critical subspaces. Regarding the two points discussed above, Table LABEL:appendax7 and LABEL:appendax8 provide detailed empirical evidence. Both On-Policy Distillation and DPO exhibit clear low-rank structures in their update patterns, demonstrating that the phenomena identified in this work are not unique to RL. Notably, On-Policy RL—combining advantages from both approaches—shows an even more pronounced low-rank property.
@@
Figure 9: Effect of On Policy Distillation and DPO on the embedding layer, the two representations of the same token are connected with gray lines.
@@
Appendix D External Manifestations of Rank-1 Dominance
@@
First, let’s make sure I understand all the components: Triangle ABCABC: A triangle with vertices AA, $B$, and $C$. Medians ADAD, BEBE, and CFCF: A median of a triangle is a line segment joining a vertex to the midpoint of the opposite side. So: $D$ is the midpoint of $BC$. $E$ is the midpoint of $AC$. $F$ is the midpoint of $AB$. Centroid $G$: The point where the three medians intersect. It’s also the “center of mass” of the triangle. Line through $G$ parallel to BCBC: This line is parallel to side BCBC and passes through the centroid $G$. Points $M$ and NN: The intersections of this parallel line with sides ABAB and $AC$, respectively. Area of triangle ABCABC: Given as 144. Goal: Find the area of triangle $ENG$. Note that point $E$ is already defined as the midpoint of ACAC, so triangle ENGENG has vertices $E$, NN, and $G$.
@@
Appendix E Linear Projection Methods